OpenStudy (anonymous):

Find the values of x where the extreme values of the function y=(x^3)-7x-5x occur. A. x=-0.33, x=5 B. x=0.33, x=-5 C. x=0.33, x=5 D. x=2, x=5 E. x=0.65, x=7.65 show me how you got the answer Please

OpenStudy (anonymous):

is that function right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay well then we can simplify it first by collecting like term: y=(x^3)-7x-5x y=x^3-12x then we find the derivative y'=3x^2-12 next we set the derivative to zero: 3x^2-12=0-->3(x^2-4)=0-->3(x-2)(x+2)=0 Thus our critical number are: x-2=0--->x=2 and x+2=0--->x=-2 Which is of course makes no sense since this not one of your possible choices

OpenStudy (anonymous):

maybe we shouldn't combine like terms in the beginning .

OpenStudy (anonymous):

even if you dont we will still get the same answer: y=x^3-7x-5x y'=3x^2-7-5 =3x^2-12, see same thing

OpenStudy (anonymous):

i suggest you check if that is the correct function

OpenStudy (anonymous):

\[y=x^3-7x-5x\]

OpenStudy (anonymous):

I'm not lying to you, this function is never gonna have extreme point on one of the choices you posted: http://www.wolframalpha.com/input/?i=y%3Dx%5E3-7x-5x

OpenStudy (anonymous):

by extreme points you mean critical points right? max and min

OpenStudy (anonymous):

yes